# Inexact Moreau Envelope Lagrangian Method for Non-Convex Constrained Optimization under Local Error Bound Conditions on Constraint Functions

### v1

@@ 5 Convergence Analysis for the iMELa Method
Suppose that Assumptions 3.1 and 3.3 hold and, in Algorithm 1,
$\tau_{t}=\tau=\frac{p-L}{4B_{g}^{2}}\text{ and }\theta_{t}=\theta=\min\left\{\frac{p-L}{18p},6p\tau,\frac{(p-L)\tau}{12p\gamma K},\frac{R^{2}(\delta)(p-L)^{2}\tau}{2\cdot 12^{3}p^{3}M_{\boldsymbol{\lambda}}^{2}}\right\},\forall\,t\geq 0.$
$\phi^{t}-\phi^{t+1}\geq$ $\leavevmode\nobreak\ \frac{p-L}{4}\|{\mathbf{x}}^{(t)}-{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})\|^{2}+\frac{p}{12\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}$
$+\frac{\tau}{2}\textup{dist}^{2}\left({\mathbf{g}}({\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\right)-\frac{p\theta\epsilon_{t}^{2}}{12(p-L)^{2}}.$ (68)
Proof.
By the choices of $\tau_{t}$ and $\theta_{t}$, inequality (41) holds by Proposition 5.5. We then prove (68) separately for the cases when (51) is satisfied and when it is not.
Case I: (51) is satisfied. Inequality (54) holds by Proposition 5.7. Hence,
$\leavevmode\nobreak\ \tau\cdot\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)},\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
(${\geq}$ $\leavevmode\nobreak\ \tau\cdot\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)},\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}-6p\theta\frac{\gamma K}{p-L}\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)},\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}$
[…]
where the second inequality is because $\theta\leq\frac{(p-L)\tau}{12p\gamma K}$. Then applying (69) to (41) implies (68).
Case II: condition (51) is not satisfied. In this case, it holds that
$\text{dist}^{2}\big{(}-\nabla f(\tilde{{\mathbf{x}}}^{(t+1)})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\boldsymbol{\lambda}^{(t+1)},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)})\big{)}$ (70)
$+\;\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}$ $>R^{2}(\delta).$
By the triangle inequality and $-\nabla_{\mathbf{x}}\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\in\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)})$, it holds that
$\leavevmode\nobreak\ \text{dist}^{2}\big{(}-\nabla f(\tilde{{\mathbf{x}}}^{(t+1)})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\boldsymbol{\lambda}^{(t+1)},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)})\big{)}$
$\leq$ $\leavevmode\nobreak\ p^{2}\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\leq 2p^{2}\big{(}\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\|^{2}+\|{\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\big{)}$
$=$ $\leavevmode\nobreak\ 2p^{2}\big{(}\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\|^{2}+(1/\theta^{2})\cdot\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\big{)}$
$\leq$ $\leavevmode\nobreak\ 2p^{2}\left(\frac{\epsilon_{t}^{2}}{(p-L)^{2}}+\frac{1}{\theta^{2}}\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\right),$ (71)
$\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}+2p^{2}\left(\frac{\epsilon_{t}^{2}}{(p-L)^{2}}+\frac{1}{\theta^{2}}\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\right)>R^{2}(\delta).$ (72)
We then have
$\leavevmode\nobreak\ \frac{p}{6\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\tau\cdot\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
(${\geq}$ $\leavevmode\nobreak\ \frac{p}{6\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\tau\cdot\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}$
$-\frac{12p\theta M_{\boldsymbol{\lambda}}}{p-L}\text{dist}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}$
[…]
$+\frac{R^{2}(\delta)\theta}{24p}-\frac{p\theta\epsilon_{t}^{2}}{12(p-L)^{2}}$
[…]
Building on Proposition 5.9, we establish the complexity of Algorithm 1 in the following theorem.
@@
Suppose that the assumptions in Proposition 5.9 hold. If $\epsilon_{t}=\frac{c}{t+1}$ for any constant $c>0$ and
@@
$T=\left[\left(1+\frac{(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\frac{c\pi^{2}}{3}+\max\left\{\frac{24p}{\theta},\frac{4(M_{\boldsymbol{\lambda}}^{2}+1)}{\tau}\right\}\left(\phi^{0}-\underline{f}+\frac{cp\theta\pi^{2}}{72(p-L)^{2}}\right)\right]\Big{/}\epsilon^{2},$
Algorithm 1 finds an $\epsilon$-KKT solution for (1) with oracle complexity $O(\epsilon^{-2}\ln(\epsilon^{-1}))$.
Proof.
$\leavevmode\nobreak\ \text{dist}\big{(}-\nabla f({\mathbf{x}}^{(t+1)})-\textstyle\sum_{i=1}^{m}\lambda_{i}^{(t+1)}\cdot\nabla g_{i}({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\big{)}$
$\leq$ $\leavevmode\nobreak\ \text{dist}\big{(}-\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\big{)}+p\|{\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)}\|$
$\leq$ $\leavevmode\nobreak\ \epsilon_{t}+p\|{\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)}\|=\epsilon_{t}+(p/\theta)\cdot\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|.$ (74)
$\|[{\mathbf{g}}({\mathbf{x}}^{(t+1)})]_{+}\|\leq$ $\leavevmode\nobreak\ \|[{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})]_{+}\|+\|[{\mathbf{g}}({\mathbf{x}}^{(t+1)})]_{+}-[{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})]_{+}\|$
$\leq$ $\leavevmode\nobreak\ \|[{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})]_{+}\|+B_{g}\|{\mathbf{x}}^{(t+1)}-\tilde{{\mathbf{x}}}^{(t+1)}\|$
$\leq$ $\leavevmode\nobreak\ \|[{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})]_{+}\|+\frac{B_{g}\epsilon_{t}}{p-L}\leq\text{dist}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}+\frac{B_{g}\epsilon_{t}}{p-L},$ (75)
$\textstyle\sum_{i=1}^{m}|\lambda_{i}^{(t+1)}g_{i}({\mathbf{x}}^{(t+1)})|=$ $\leavevmode\nobreak\ \textstyle\sum_{i\in I_{g}^{(t+1)}}|\lambda_{i}^{(t+1)}g_{i}({\mathbf{x}}^{(t+1)})|$
$\leq$ $\leavevmode\nobreak\ \textstyle\sqrt{\sum_{i\in I_{g}^{(t+1)}}(\lambda_{i}^{(t+1)})^{2}}\sqrt{\sum_{i\in I_{g}^{(t+1)}}g^{2}_{i}({\mathbf{x}}^{(t+1)})}$
$\leq$ $\leavevmode\nobreak\ M_{\boldsymbol{\lambda}}\cdot\text{dist}\big{(}{\mathbf{g}}({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}$
$\leq$ $\leavevmode\nobreak\ M_{\boldsymbol{\lambda}}\cdot\text{dist}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}+M_{\boldsymbol{\lambda}}\cdot\|{\mathbf{g}}({\mathbf{x}}^{(t+1)})-{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\|$
$\leq$ $\leavevmode\nobreak\ M_{\boldsymbol{\lambda}}\cdot\text{dist}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}+\frac{M_{\boldsymbol{\lambda}}B_{g}\epsilon_{t}}{p-L},$ (76)
where the first inequality is by the Cauchy-Schwarz inequality, the second one is by Lemma 4.1 and the fact that $g^{2}_{i}({\mathbf{x}}^{(t+1)})=\text{dist}^{2}(g_{i}({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}}(\lambda_{i}^{(t+1)}))$ for $i\in I_{g}^{(t+1)}$, the third one is by the triangle inequality, and the last one is by (22), (56) with ${\mathbf{x}}={\mathbf{x}}^{(t+1)}$, and Assumption 3.1E.
By summing up (68) for $t=0,1,\dots,T-1$, we have
$\leavevmode\nobreak\ \sum_{t=0}^{T-1}\left(\frac{p-L}{4}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}+\frac{p}{12\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\frac{\tau}{2}\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}\right)$
$\leq$ $\leavevmode\nobreak\ \phi^{0}-\phi^{T}+\frac{p\theta}{12(p-L)^{2}}\sum_{t=0}^{T-1}\epsilon_{t}^{2}\leq\phi^{0}-\underline{f}+\frac{cp\theta\pi^{2}}{72(p-L)^{2}},$ (77)
where the second inequality is by (34) and the fact that
$\sum_{t=0}^{T-1}\epsilon_{t}^{2}\leq\sum_{t=0}^{\infty}\epsilon_{t}^{2}=\frac{c\pi^{2}}{6}.$ (78)
$\leavevmode\nobreak\ \sum_{t=0}^{T-1}\left(\begin{array}[]{l}\text{dist}^{2}\big{(}-\nabla f({\mathbf{x}}^{(t+1)})-\sum_{i=1}^{m}\lambda_{i}^{(t+1)}\cdot\nabla g_{i}({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\big{)}\\ +\|[{\mathbf{g}}({\mathbf{x}}^{(t+1)})]_{+}\|^{2}+\big{(}\sum_{i=1}^{m}|\lambda_{i}^{(t+1)}g_{i}({\mathbf{x}}^{(t+1)})|\big{)}^{2}\end{array}\right)$ (81)
$\leq$ $\leavevmode\nobreak\ \frac{2p^{2}}{\theta^{2}}\sum_{t=0}^{T-1}\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}+\left(2+\frac{2(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\sum_{t=0}^{T-1}\epsilon_{t}^{2}$
$+2(M_{\boldsymbol{\lambda}}^{2}+1)\cdot\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}$
$\leq$ $\leavevmode\nobreak\ \left(1+\frac{(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\frac{c\pi^{2}}{3}+\max\left\{\frac{24p}{\theta},\frac{4(M_{\boldsymbol{\lambda}}^{2}+1)}{\tau}\right\}\left(\phi^{0}-\underline{f}+\frac{cp\theta\pi^{2}}{72(p-L)^{2}}\right),$ (82)
where the second inequality is because of (77) and (78). According to the choice of $T$, there exists $s\in\{0,1,\dots,T-1\}$ such that
$\leavevmode\nobreak\ \text{dist}^{2}\big{(}-\nabla f({\mathbf{x}}^{(s+1)})-\textstyle\sum_{i=1}^{m}\lambda_{i}^{(s+1)}\cdot\nabla g_{i}({\mathbf{x}}^{(s+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(s+1)})\big{)}$ (83)
$+\|[{\mathbf{g}}({\mathbf{x}}^{(s+1)})]_{+}\|^{2}+\big{(}\textstyle\sum_{i=1}^{m}|\lambda_{i}^{(s+1)}g_{i}({\mathbf{x}}^{(s+1)})|\big{)}^{2}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \epsilon^{2}.$
According to Definition 3.2, ${\mathbf{x}}^{(s+1)}$ is an $\epsilon$-KKT point of (1). By (31), the complexity of the $t$-th iteration of Algorithm 1 is $O\left(\sqrt{\frac{K}{p-L}}\ln(\epsilon_{t}^{-1})\right)=O(\ln(\epsilon^{-1}))$ for $t=0,1,\dots,T-1$. Therefore, the total oracle complexity of Algorithm 1 for finding an $\epsilon$-KKT point is $O(\epsilon^{-2}\ln(\epsilon^{-1}))$. ∎
@@ 6 Numerical Experiment
We demonstrate the performance of the iMELa method on a fairness-aware classification problem, which is an instance of (1). We compare it with the inexact proximal point penalty (iPPP) method [28], the smoothed proximal Lagrangian method (SP-LM) [39], and the switching subgradient (SSG) method [16]. All experiments were conducted using MATLAB 2024a on a computer with an Intel Core i7-7700 CPU (3.60 GHz) and 32 GB of RAM.
Recall that, at its $t$-th iteration, the iMELa method obtains
where $\rho_{t}>0$ is the penalty parameter and $p_{t}>0$ is the proximal parameter. For a fair comparison, we apply the accelerated projected gradient (APG) method in [36, (2.2.22)] to approximately solve (84) and (85). Let $F({\mathbf{u}})$ be the objective function in (84) or (85) and let ${\mathbf{u}}^{(k)}$ be the main iterate generated by the APG method. The APG method is terminated when
where $T_{1/\eta_{t}}({\mathbf{u}}^{(k)}):=\text{proj}_{\mathcal{X}}\left({\mathbf{u}}^{(k)}-\eta_{t}\nabla F({\mathbf{u}}^{(k)})\right)$, $\eta_{t}$ is a step-size and $\epsilon_{t}^{\prime}$ is a targeted tolerance. Then we take ${\mathbf{x}}^{(t+1)}=T_{1/\eta_{t}}({\mathbf{u}}^{(k)})$ as the approximate solution to (84) or (85). Here, $\mathcal{G}_{1/\eta_{t}}({\mathbf{u}}^{(k)})$ is known as the gradient mapping of $F(\cdot)$ at ${\mathbf{u}}^{(k)}$. By [35, (2.15)], (86) ensures $\text{dist}\left(-\nabla F({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\right)\leq\epsilon_{t}$ with some $\epsilon_{t}=O(\epsilon_{t}^{\prime})$, i.e., condition (19) holds if $F$ is from (84).
In addition, at its $t$-th iteration, SP-LM updates ${\mathbf{x}}$ by
and, at its $t$-th iteration, the SSG method updates ${\mathbf{x}}$ by
${\mathbf{x}}^{(t+1)}=\begin{cases}\text{proj}_{\mathcal{X}}({\mathbf{x}}^{(t)}-\eta_{t}\nabla f({\mathbf{x}}^{(t)}))\text{ if }\max_{i\in[m]}g_{i}({\mathbf{x}}^{(t)})\leq\epsilon_{t},\\ \text{proj}_{\mathcal{X}}({\mathbf{x}}^{(t)}-\eta_{t}\boldsymbol{\zeta}_{g}^{(t)})\text{ for some }\boldsymbol{\zeta}_{g}^{(t)}\in\mathrm{conv}(\{\nabla g_{j}({\mathbf{x}}^{(t)}):j\in I({\mathbf{x}}^{(t)})\})\text{ otherwise,}\end{cases}$ (88)
Next, we define the instance of (1) solved in our experiments. Given a feature vector ${\mathbf{a}}\in\mathbb{R}^{d}$ and a class label $b\in\{1,-1\}$, the goal of linear binary classification is to learn a model ${\mathbf{x}}\in\mathbb{R}^{d}$ to predict $b$ based on the score ${\mathbf{x}}^{\top}{\mathbf{a}}$ with a larger score indicating a higher chance of a positive label. Suppose that a training set $\mathcal{D}=\{({\mathbf{a}}_{i},b_{i})\}_{i=1}^{n}$ is available. A model ${\mathbf{x}}$ can be obtained by solving
@@
where $\kappa$ is the slackness parameter indicating how much we are willing to increase the loss in order to reduce $|\mathcal{R}({\mathbf{x}})|$ to obtain a more fair model, and $\mathcal{X}$ is the same as that in (89). In the experiments, we set $\mathcal{X}=\{{\mathbf{x}}\in\mathbb{R}^{d}:\|{\mathbf{x}}\|_{1}\leq r\}$ to satisfy Assumption 3.1A.
We solve problem (91) on three datasets: a9a [18], bank [34] and COMPAS [1]. Details about these datasets are provided in Table 1. Each dataset is splited into two subsets with a ratio of 2 : 1. The larger subset serves as $\mathcal{D}$ in the constraint, while the smaller subset is further partitioned into $\mathcal{D}_{p}$ and $\mathcal{D}_{u}$ based on the binary group variable specified in Table 1.
Datasets $n$ $d$ Label Groups
a9a 48,842 123 Income Gender
Bank 41,188 54 Subscription Age
COMPAS 6,172 16 Recidivism Race
The methods in comparison are initialized at ${\mathbf{x}}^{(0)}={\mathbf{x}}_{\text{feas}}$ obtained above. Based on the definitions of $\mathcal{R}$ and $\mathcal{L}$, we can calculate an upper bound of the smooth parameter of the objective and constraint functions in (91) using the data. For the objective function in (91), we notice that $\mathcal{R}({\mathbf{x}})\in[0,1]$ for any ${\mathbf{x}}\in\mathbb{R}^{d}$ and, by a derivation similar to [16, (75-77)], $\mathcal{R}({\mathbf{x}})$ and $\nabla\mathcal{R}({\mathbf{x}})$ are Lipschitz continuous with constants of
respectively. Therefore, (91) satisfies Assumption 3.1 with any $L\geq\max\{\beta+\alpha^{2},\gamma^{\prime}\}$. After estimating $\max\{\beta+\alpha^{2},\gamma^{\prime}\}$ using data, we set $L$ to 10, 10 and 2.5 on a9a, bank and COMPAS, respectively, and set $p=2L$. In the APG method applied to (84) and (85), we use a constant step-size $\eta$. On a9a and bank, we select $\eta$ from $\{0.005,0.01,0.02,0.05\}$; on COMPAS, we select $\eta$ from $\{0.02,0.05,0.1,0.2\}$. The choices of other parameters are uniform across all datasets. For the iMEla method, we select $\tau_{t}=\tau$ from $\{5,10,20,50\}$ and $\theta_{t}=\theta$ from $\{0.5,0.75,1\}$, and set $\epsilon_{t}=\frac{c}{t+1}$ with $c$ selected from $\{1,2,5,10\}$. For the iPPP method, following [28], we set $\rho_{t}=\rho\sqrt{t+1}$ and $\epsilon_{t}=\frac{1}{\rho_{t}(t+1)}$ with $\rho$ selected from $\{200,500,1000,1500\}$. For SP-LM, we select $\eta$ from $\{0.005,0.01,0.02,0.05\}$ on a9a and bank and from $\{0.02,0.05,0.1,0.2\}$ on COMPAS. We then select $\tau_{t}=\tau$ from $\{5,10,20,50\}$ and $\theta_{t}=\theta$ from $\{0.5,0.75,1\}$. For the SSG method, following [16], we adopt both static and diminishing step-sizes. For the static step-size, we select $\epsilon_{t}=\epsilon$ from $\{10^{-6},2\times 10^{-6},5\times 10^{-6},10^{-5}\}$ and $\eta_{t}=\eta$ from $\{2\times 10^{-4},5\times 10^{-4},10^{-3},2\times 10^{-3}\}$. For the diminishing step-size, we select $\epsilon_{t}=\frac{E_{1}}{\sqrt{t+1}}$ and $\eta_{t}=\frac{E_{2}}{\sqrt{t+1}}$ and select $E_{1}$ from $\{5\times 10^{-5},10^{-4},2\times 10^{-4},5\times 10^{-4}\}$ and $E_{2}$ from $\{0.02,0.05,0.1,0.2\}$. Since the iMELa method, the iPPP method and SP-LM are primal-dual based, their best combination of parameters is chosen as the one that minimizes the smallest value of
$\leavevmode\nobreak\ \text{dist}\big{(}-\mathcal{R}({\mathbf{x}}^{(s)})\cdot\nabla\mathcal{R}({\mathbf{x}}^{(s)})-\lambda^{(s)}\cdot\nabla\mathcal{L}({\mathbf{x}}^{(s)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(s)})\big{)}$
among iterations $s=0,1,\dots,T_{\text{tuning}}-1$. Here, $T_{\text{tuning}}$ is the number of iterations for parameter tuning. Note that $\boldsymbol{\lambda}^{(t)}$ is directly generated in the iMELa method and SP-LM, but we need to set $\boldsymbol{\lambda}^{(t)}=\rho_{t-1}[{\mathbf{g}}({\mathbf{x}}^{(t)})]_{+}$ for the iPPP method according to Step 5 of Algorithm 1 in [28]. Since the SSG method does not generate $\boldsymbol{\lambda}^{(t)}$, its best combination of parameters is identified as the one that minimizes the smallest value of $\frac{1}{2}(\mathcal{R}({\mathbf{x}}^{(s)}))^{2}$ when $\mathcal{L}({\mathbf{x}}^{(s)})\leq\mathcal{L}^{*}+\kappa+10^{-5}$ among iterations $s=0,1,\dots,T_{\text{tuning}}-1$. For SP-LM and the SSG method, we set $T_{\text{tuning}}=1000$, 5000 and 75000 for a9a, bank and COMPAS, respectively. For a fair comparison, for the iMEla method and the iPPP method, we set $T_{\text{tuning}}$ to be the smallest integer satisfying $\sum_{t=0}^{T_{\text{tuning}}-1}k_{t}\geq 1000\text{, }5000\text{ and }75000$ for a9a, bank and COMPAS, respectively, where $k_{t}$ is the number of inner steps performed at the $t$-th outer iteration. Once all parameters are chosen, we run SP-LM and the SSG method for $T=4T_{\text{tuning}}$ iterations, and run the iMEla method and the iPPP method for $T$ iterations with $T$ being the smallest integer satisfying $\sum_{t=0}^{T-1}k_{t}\geq 4\sum_{t=0}^{T_{\text{tuning}}-1}k_{t}$.

### v2

@@ 5 Convergence Analysis for the iMELa Method
Suppose that Assumptions 3.1 and 3.3 hold and, in Algorithm 1, we set
$\begin{aligned} \tau_{t}=\tau=&\penalty 10000\ \frac{p-L}{4B_{g}^{2}},\\ \theta_{t}=\theta\leq&\penalty 10000\ \tilde{\theta}_{d}:=\min\left\{\frac{p-L}{18p},6p\tau,\frac{R^{2}(\delta)(p-L)^{2}\tau}{2\cdot 12^{3}p^{3}M_{\boldsymbol{\lambda}}^{2}},\hat{\theta}_{d}\right\},\end{aligned}$ (34)
for $\forall\,t\geq 0$, where $C_{d}:=\frac{\gamma^{1/d}K}{p-L}$ and
$\hat{\theta}_{d}:=\frac{\tau}{12pC_{d}}\left(\frac{\tau}{12p}\right)^{1-1/d}.$ (35)
Then the sequence $\{({\mathbf{x}}^{(t)},\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})\}_{t\geq 0}$ generated by Algorithm 1 satisfies that
$\phi^{t}-\phi^{t+1}\geq$ $\penalty 10000\ \frac{p-L}{4}\|{\mathbf{x}}^{(t)}-{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})\|^{2}+\frac{p}{12\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}$
$+\frac{\tau}{2}\textup{dist}\big({\mathbf{g}}({\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-\Xi_{d}(\theta)-\frac{p\theta\epsilon_{t}^{2}}{12(p-L)^{2}},$ (36)
where
$\Xi_{d}(\theta):=\begin{cases}\left(1-\dfrac{1}{d}\right)\left(\dfrac{2}{\tau d}\right)^{\frac{1}{d-1}}\big(6p\theta\,C_{d}\big)^{\frac{d}{d-1}},&d\in(1,2],\\[4.30554pt] 0,&d=1.\end{cases}$ (37)
By Proposition 5.5, we establish the complexity of Algorithm 1 in the following theorem whose proof is in Section D.
@@
Suppose that the assumptions in Proposition 5.5 hold with a fixed exponent $d\in[1,2]$. For a given target accuracy $\epsilon>0$, choose $\theta_{t}=\theta=\theta(\epsilon)$, where
@@
$\theta(\epsilon):=\min\left\{\tilde{\theta}_{d},\,\frac{6p\tau}{M_{\boldsymbol{\lambda}}^{2}+1},\,\left(\frac{\epsilon^{2}}{2A_{d}}\right)^{d-1}\right\},\penalty 10000\ \text{ with }\penalty 10000\ A_{d}:=24p\left(1-\frac{1}{d}\right)\left(\frac{2}{\tau d}\right)^{\frac{1}{d-1}}(6pC_{d})^{\frac{d}{d-1}},$
and $\tilde{\theta}_{d}$, $C_{d}$ (and $\tau$) defined in Proposition 5.5. If $\epsilon_{t}=\frac{c}{t+1}$ for any constant $c>0$ and
$\begin{aligned} T:=\left\lceil\frac{2\mathcal{C}_{0}(\theta(\epsilon))}{\epsilon^{2}}\right\rceil,\text{ where }\penalty 10000\ \mathcal{C}_{0}(\theta):=&\penalty 10000\ \left(\begin{array}[]{l}\left(1+\frac{(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\frac{c\pi^{2}}{3}\\ +\;\mathcal{C}_{1}(\theta)\left(\phi^{0}-\underline{f}+\frac{cp\theta\pi^{2}}{72(p-L)^{2}}\right)\end{array}\right),\\ \mathcal{C}_{1}(\theta):=&\penalty 10000\ \max\left\{\frac{24p}{\theta},\frac{4(M_{\boldsymbol{\lambda}}^{2}+1)}{\tau}\right\},\end{aligned}$ (38)
Algorithm 1 finds an $\epsilon$-KKT solution for (1) with oracle complexity $O(\epsilon^{-2d}\ln(\epsilon^{-1}))$.
@@ 6 Numerical Experiment
We demonstrate the performance of the iMELa method on a fairness-aware classification problem, which is an instance of (1). We compare it with the inexact proximal point penalty (iPPP) method [28], the damped proximal augmented Lagrangian method (DPALM) [8],the smoothed proximal Lagrangian method (SP-LM) [40], and the switching subgradient (SSG) method [16].
First, we define the instance of (1) solved in our experiments. Given a feature vector ${\mathbf{a}}\in\mathbb{R}^{d}$ and a class label $b\in\{1,-1\}$, the goal of linear binary classification is to learn a model ${\mathbf{x}}\in\mathbb{R}^{d}$ to predict $b$ based on the score ${\mathbf{x}}^{\top}{\mathbf{a}}$ with a larger score indicating a higher chance of a positive label. Suppose that a training set $\mathcal{D}=\{({\mathbf{a}}_{i},b_{i})\}_{i=1}^{n}$ is available. A model $\mathbf{x}$} can be obtained by solving